# Generative Psycho-Lexical Approach for Constructing Value Systems in Large Language Models

### v5

@@ 1 Introduction
However, value systems used in existing research—primarily Schwartz’s values—are established for humans. A psychologically informed value system tailored for LLMs remains largely unexplored, obscuring a structured and holistic perspective on their values. A preliminary attempt at constructing LLM value systems Biedma et al. (2024) has several limitations that undermine its psychological grounding; we present related discussion and experimental evidence in § 2.3 and Appendix C.
@@
First, GPLA supports agent-specific construction. Rather than relying on value lexicons derived from human-compiled dictionaries, it utilizes LLMs to identify values within free-form agent-generated text. This principled approach facilitates the agent- and context-specific collection of value lexicons and ensures better coverage (§ B.1). Second, GPLA enables nonreactive value measurement. Traditional methods that rely on self-report are demonstrated to be unreliable and biased for humans Ponizovskiy et al. (2020) and AI Dominguez-Olmedo et al. (2023); Röttger et al. (2024); Ye et al. (2025b). Instead, GPLA measures values within unstructured responses to arbitrary prompts, which mitigates response biases and enhances scalability and validity Ye et al. (2025b).
@@ 2.3 From Human Values to LLM Values
One preliminary attempt Biedma et al. (2024) at constructing LLM value systems faces several limitations that undermine its psychological grounding: 1) the value lexicon collection is susceptible to response bias and lacks comprehensive coverage; 2) the derivation of value correlations lacks theoretical or empirical validity; and 3) the resulting value system has yet to be evaluated in terms of its reliability, validity, or utility. Appendix C presents further discussion and experimental evidence.
@@ 3 Generative Psycho-Lexical Approach for Constructing LLM Value System
Generative Psycho-Lexical Approach (GPLA), as our foundation for constructing the LLM value system, is illustrated schematically in Fig. 1 and algorithmically in Algorithm 1. GPLA adopts an agentic framework, comprising three LLM agents and four sequential steps, to systematically collect, filter, measure, and structure the LLM value system.
@@
3:  $P=M_{P}(\mathcal{C})$ // Extract perceptions $P$ from corpus $\mathcal{C}$
4:  $\{V,Freq\}=M_{G}(P)$ // Generate values $V$ behind perceptions $P$ and collect their frequencies
5:  $V=\text{LexiconFiltering}(V,Freq)$ // Remove duplicate and less prioritized values
6:  $\mathbf{X}_{M}=\text{GPV}(V,M;M_{P},M_{E})$ // Measure the value orientations of LLM subjects
7:  $V_{H},\mathbf{\Lambda}=\text{PCA}(\mathbf{X}_{M})$ // Compute the hidden factors and structure of the system
@@
In line with the established psycho-lexical approach De Raad et al. (2016), we begin by comprehensively identifying atomic values that LLMs may endorse. We draw upon a diverse, value-laden corpus of LLM outputs $\mathcal{C}$ from ValueBench Ren et al. (2024), GPV Ye et al. (2025b), BeaverTails Ji et al. (2024), and ValueLex Biedma et al. (2024). These corpora provide a broad spectrum of value-laden LLM generations. Detailed statistics and further explanations on these datasets are provided in § A.1.
@@
Then, we extract perceptions $P$ from the collected corpus using the Perception Parser $M_{P}$. Perceptions are atomic value-rich expressions that reflect values behind free-form LLM outputs and are akin to stimuli in traditional psychometric tools Ye et al. (2025b). These perceptions are then mapped to underlying values $V$ through the Value Generator $M_{G}$, with their frequency of occurrence $Freq$ recorded. We instantiate $M_{P}$ following GPV Ye et al. (2025b) and $M_{G}$ using Kaleido model Sorensen et al. (2024a). In § B.1, we confirm the comprehensive coverage of the generated values.
@@
Structuring the LLM value system requires identifying correlations between values based on LLM value measurements. To this end, we adopt GPV Ye et al. (2025b), instantiated with $M_{P}$ and $M_{E}$, to measure the value orientations $\mathbf{X}_{M}$ of LLM subjects $M$ on values $V$. GPV dynamically extracts psychometric stimuli from LLM outputs and reveals value orientations therein (detailed in § A.3). We measure a set of 693 LLM subjects, consisting of 33 LLMs paired with 21 profiling prompts (listed in § A.2). The involved LLMs are instruction-tuned models, as they are more relevant for real-world, public-facing applications. The value system $\mathcal{VS}=(V,V_{H},\mathbf{\Lambda})$ is then constructed through Principal Component Analysis (PCA) Ponizovskiy et al. (2020), which uncovers the underlying value factors $V_{H}$ and hierarchical relationships $\mathbf{\Lambda}$ (loadings of $V$ on $V_{H}$).
@@ 4.3 LLM Value Alignment for Evaluating Representation Power
Figure 2: Dendrogram of our value system. Values with a "*" are negatively loaded.
Figure 3: Correlation heatmap and circumplex analysis.
@@ 5.1 Proposed Value System
Fig. 3 visualizes the value system constructed by GPLA. Table 1 gathers its partial factor loadings, Cronbach’s Alpha ($\alpha$), and confidence intervals (CI). The factor loadings indicate the strength of the relationship between each factor and its atomic values, with higher loadings suggesting more contribution to the factor. Cronbach’s Alpha measures the internal consistency of each factor, with higher values indicating greater reliability. Some atomic values are removed to ensure clear loading patterns and desirable factor reliability Aavik and Allik (2002). After that, all our factors reach the standard psychometric threshold of 0.7, indicating strong internal consistency Taber (2018).
Analyzing factor loadings reveals the underlying structure of five key factors: Social Responsibility emphasizes fairness, inclusivity, and societal well-being; Risk-Taking reflects openness to change and embracing uncertainty; Rule-Following highlights order, discipline, and pragmatism; Self-Competence focuses on personal growth and achievement; and Rationality prioritizes logical, evidence-based decision-making grounded in objectivity and impartiality. Full factor loadings and further analysis are available in § B.3.
@@ 5.2 Analyzing Value System
Fig. 3 (Top) presents the correlations between the factors in our value system. Similar to Schwartz’s theory of basic human values Schwartz (2012), LLM values also exhibit both compatible and opposing relationships. Notably, social responsibility, rule-following, and rationality show positive correlations with one another, while all of them are negatively correlated with risk-taking.
@@
Circumplex analysis is a statistical method that examines whether the underlying structure of variables aligns with a circumplex pattern, and, if so, the positions of variables on a circle. The stronger the correlation between variables, the shorter their distance on the circumference. We conduct circumplex analysis based on the correlations between factors. Fig. 3 (Bottom) illustrates the analysis results based on Browne’s circular stochastic process model Browne (1992); Grassi et al. (2010). The compatible values are closer on the circle (e.g., Social Responsibility and Rule-Following) while opposing values are positioned diagonally (e.g., Risk-Taking and Rule-Following). The results verify the presence of a circumplex structure in the value system.
@@
To evaluate the consistency and robustness of our multivariate system structure (i.e., the 5-dimensional relationship), we measure LLM values using two distinct prompt datasets: the psychometric prompts from GPV Ye et al. (2025b) and the red-teaming prompts from SALAD-Bench Li et al. (2024b), which feature distinct prompt distributions. Their measurements yield an average intra-LLM correlation of 0.87; here we use intra-LLM correlations because the relative value hierarchy within an LLM is more important than their absolute measurements. This result indicates a high level of consistency in the value structure across prompt distributions. We also find a high correlation (0.73) between intra-LLM value consistency and LLM safety scores Li et al. (2024b). It suggests that LLMs with higher value consistency tend to be safer. Complete results are available in § B.4.
@@ 5.3 Comparing Value Systems
We follow the standard validation procedures of CFA Schwartz and Boehnke (2004) to evaluate the structure validity of different value systems. For our value system, we construct it using half of the measurement data and bootstrap the data to ensure its sufficiency (#data points 5$\geq$ 5$\times$ 5 × #variables). The other half of the data is held out for CFA. For Schwartz’s value systems, we map the observed value variables (the atomic values in our system) to its four high-level values or ten low-level values, according to the semantic relevance Schwartz and Boehnke (2004) measured by an embedding model OpenAI (2024); all data is used for CFA. Table 2 displays the CFA results. Our value system demonstrates a better fit for the data, capturing the underlying values behind LLM generations.
@@
We follow the setup in (Ye et al., 2025b, Section 5.2) to predict LLM safety based on their value orientations. Table 3 presents the prediction accuracy under different value systems. The higher accuracy of our value system indicates its superior predictive validity. In addition, according to the parameters of well-trained linear classifiers, we find that Social Responsibility, Rule-Following, and Rationality enhance safety, whereas Risk-taking and Self-Competence undermine it; see § B.6 for full results.
@@
We follow the experimental setup in (Yao et al., 2024a, Section 6.2) to perform LLM value alignment. Different value systems are respectively used to represent LLM outputs and desired human values. We employ GPV Ye et al. (2025b) as an open-vocabulary value evaluator for all value systems, but also include the original results of Yao et al. (2024a) using a Schwartz-specific evaluator for comparison. Table 4 shows the alignment performance of different value systems. Alignment under our value system converges to the lowest harmlessness and the highest helpfulness, establishing its superior representation power. Full experimental details are available in § A.4.
@@ 6.2 Advantages of GPLA over Prior Psycho-Lexical Approaches
Here we discuss the main advantages of GPLA over the traditional psycho-lexical approach in constructing value systems. We defer the comparison with ValueLex Biedma et al. (2024), a recent study that constructs LLM value systems, to Appendix C.
@@
  • Comprehensive Coverage. LLMs can generate values in diverse linguistic forms, beyond words in dictionaries and thesauruses, and thus provide more comprehensive coverage of values. Please refer to § B.1 for the empirical evidence.
@@ Acknowledgements
This work is supported by the National Natural Science Foundation of China (Grant No. 62276006); Wuhan East Lake High-Tech Development Zone National Comprehensive Experimental Base for Governance of Intelligent Society; and the Fundamental Research Funds for the Central Universities.
@@ A.1 Data Statistics for Collecting Value Lexicons
We collect value-laden LLM generations from four data sources: ValueBench Ren et al. (2024), GPV Ye et al. (2025b), ValueLex Biedma et al. (2024), and BeaverTails Ji et al. (2024). They provide data of different forms: raw LLM responses, parsed perceptions (a sentence that is highly reflective of values Ye et al. (2025b)), and annotated values. The summary of the data statistics is shown in Table 5.
@@ A.2 LLM Subjects
Our experiments involve 33 LLMs (Table 6) and 21 profiling prompts, including the default prompt (’You are a helpful assistant’) Zhu et al. (2023) and 20 value-anchoring prompts Rozen et al. (2024).
@@ A.4 LLM Value Alignment
All experiments were conducted on two NVIDIA L20 GPUs, each with 48GB of memory. We generally follow the experimental setup in Yao et al. (2024a), with the exceptions noted below. As shown in Table 4, our modifications improve the harmlessness of the aligned model with only marginal reduction in helpfulness.
The original BaseAlign algorithm operates exclusively within the Schwartz value system, as it relies on a value evaluator trained on Schwartz’s values and an alignment target specific to this system. We extend BaseAlign to align LLMs under any arbitrary value system. First, we employ GPV Ye et al. (2025b) as an open-vocabulary value evaluator. Second, we propose a method for distilling the alignment target from human preference data (§ 4.3). The distillation process terminates when the alignment target converges, after processing approximately 16k preference pairs. The results of this distillation are shown in Table 7 for Schwartz’s values and Table 8 for our system. The distilled alignment target, based on Schwartz’s values, closely matches the heuristically defined target in BaseAlign Yao et al. (2024a), demonstrating the effectiveness of our approach.
@@ B.2 Principal Component Analysis
We apply principal component analysis (PCA) to identify the underlying factors in our value system. The number of factors to retain is determined using both the scree plot and Cronbach’s alpha. The scree plot, shown in Fig. 4, reveals an elbow at the fifth or sixth component. We retain five factors due to their high reliability.
@@ B.4 Consistency Across Datasets
In § 5, we use the default profiling prompt for evaluating cross-dataset consistency. Table 9 displays the intra-LLM correlation of the two value measurements for each LLM. In addition, we include the model safety scores from SALAD-Bench Li et al. (2024b) for reference. The results show that the intra-LLM correlation is generally high (0.87 on average), indicating the robustness of our value system. We also find that the intra-LLM value consistency is positively correlated with its safety (0.73).
@@ B.5 Correlation with Schwartz’s Values
We also measure the Schwartz’s values Schwartz (1992) of the LLMs subjects. Table 10 presents the correlation between our factors and Schwartz’s values. The results can offer additional insights into our values.
@@ B.6 Contributions of Values to Safety
In the LLM safety prediction (§ 4.2) task, we train a linear classifier to predict the safety score of LLMs based on their values. The parameters of a well-trained classifier can be interpreted as the contributions of each value to the safety score. Such parameters, averaged over all trails, are gathered in Table 12. The results show that Social Responsibility, Rule-Following, and Rationality positively contribute to safety, while Risk-taking and Self-Competence negatively contribute to safety.

### v6

@@ 1 Introduction
However, value systems used in existing research—primarily Schwartz’s values—are established for humans. A psychologically informed value system tailored for LLMs remains largely unexplored, obscuring a structured and holistic perspective on their values. A preliminary attempt at constructing LLM value systems Biedma et al. (2024) has several limitations that undermine its psychological grounding; we present related discussion and experimental evidence in §˜2.3 and Appendix˜C.
@@
First, GPLA supports agent-specific construction. Rather than relying on value lexicons derived from human-compiled dictionaries, it utilizes LLMs to identify values within free-form agent-generated text. This principled approach facilitates the agent- and context-specific collection of value lexicons and ensures better coverage (§˜B.1). Second, GPLA enables nonreactive value measurement. Traditional methods that rely on self-report are demonstrated to be unreliable and biased for humans Ponizovskiy et al. (2020) and AI Dominguez-Olmedo et al. (2023); Röttger et al. (2024); Ye et al. (2025b). Instead, GPLA measures values within unstructured responses to arbitrary prompts, which mitigates response biases and enhances scalability and validity Ye et al. (2025b).
@@ 2.3 From Human Values to LLM Values
One preliminary attempt Biedma et al. (2024) at constructing LLM value systems faces several limitations that undermine its psychological grounding: 1) the value lexicon collection is susceptible to response bias and lacks comprehensive coverage; 2) the derivation of value correlations lacks theoretical or empirical validity; and 3) the resulting value system has yet to be evaluated in terms of its reliability, validity, or utility. Appendix˜C presents further discussion and experimental evidence.
@@ 3 Generative Psycho-Lexical Approach for Constructing LLM Value System
Generative Psycho-Lexical Approach (GPLA), as our foundation for constructing the LLM value system, is illustrated schematically in Fig.˜1 and algorithmically in Algorithm˜1. GPLA adopts an agentic framework, comprising three LLM agents and four sequential steps, to systematically collect, filter, measure, and structure the LLM value system.
@@
[…]
@@
In line with the established psycho-lexical approach De Raad et al. (2016), we begin by comprehensively identifying atomic values that LLMs may endorse. We draw upon a diverse, value-laden corpus of LLM outputs $\mathcal{C}$ from ValueBench Ren et al. (2024), GPV Ye et al. (2025b), BeaverTails Ji et al. (2024), and ValueLex Biedma et al. (2024). These corpora provide a broad spectrum of value-laden LLM generations. Detailed statistics and further explanations on these datasets are provided in §˜A.1.
@@
Then, we extract perceptions $P$ from the collected corpus using the Perception Parser $M_{P}$. Perceptions are atomic value-rich expressions that reflect values behind free-form LLM outputs and are akin to stimuli in traditional psychometric tools Ye et al. (2025b). These perceptions are then mapped to underlying values $V$ through the Value Generator $M_{G}$, with their frequency of occurrence $Freq$ recorded. We instantiate $M_{P}$ following GPV Ye et al. (2025b) and $M_{G}$ using Kaleido model Sorensen et al. (2024a). In §˜B.1, we confirm the comprehensive coverage of the generated values.
@@
Structuring the LLM value system requires identifying correlations between values based on LLM value measurements. To this end, we adopt GPV Ye et al. (2025b), instantiated with $M_{P}$ and $M_{E}$, to measure the value orientations $\mathbf{X}_{M}$ of LLM subjects $M$ on values $V$. GPV dynamically extracts psychometric stimuli from LLM outputs and reveals value orientations therein (detailed in §˜A.3). We measure a set of 693 LLM subjects, consisting of 33 LLMs paired with 21 profiling prompts (listed in §˜A.2). The involved LLMs are instruction-tuned models, as they are more relevant for real-world, public-facing applications. The value system $\mathcal{VS}=(V,V_{H},\mathbf{\Lambda})$ is then constructed through Principal Component Analysis (PCA) Ponizovskiy et al. (2020), which uncovers the underlying value factors $V_{H}$ and hierarchical relationships $\mathbf{\Lambda}$ (loadings of $V$ on $V_{H}$).
@@ 4.3 LLM Value Alignment for Evaluating Representation Power
[…]
@@ 5.1 Proposed Value System
Fig.˜7 visualizes the value system constructed by GPLA. Table˜1 gathers its partial factor loadings, Cronbach’s Alpha ($\alpha$), and confidence intervals (CI). The factor loadings indicate the strength of the relationship between each factor and its atomic values, with higher loadings suggesting more contribution to the factor. Cronbach’s Alpha measures the internal consistency of each factor, with higher values indicating greater reliability. Some atomic values are removed to ensure clear loading patterns and desirable factor reliability Aavik and Allik (2002). After that, all our factors reach the standard psychometric threshold of 0.7, indicating strong internal consistency Taber (2018).
Analyzing factor loadings reveals the underlying structure of five key factors: Social Responsibility emphasizes fairness, inclusivity, and societal well-being; Risk-Taking reflects openness to change and embracing uncertainty; Rule-Following highlights order, discipline, and pragmatism; Self-Competence focuses on personal growth and achievement; and Rationality prioritizes logical, evidence-based decision-making grounded in objectivity and impartiality. Full factor loadings and further analysis are available in §˜B.3.
@@ 5.2 Analyzing Value System
Fig.˜7 (Top) presents the correlations between the factors in our value system. Similar to Schwartz’s theory of basic human values Schwartz (2012), LLM values also exhibit both compatible and opposing relationships. Notably, social responsibility, rule-following, and rationality show positive correlations with one another, while all of them are negatively correlated with risk-taking.
@@
Circumplex analysis is a statistical method that examines whether the underlying structure of variables aligns with a circumplex pattern, and, if so, the positions of variables on a circle. The stronger the correlation between variables, the shorter their distance on the circumference. We conduct circumplex analysis based on the correlations between factors. Fig.˜7 (Bottom) illustrates the analysis results based on Browne’s circular stochastic process model Browne (1992); Grassi et al. (2010). The compatible values are closer on the circle (e.g., Social Responsibility and Rule-Following) while opposing values are positioned diagonally (e.g., Risk-Taking and Rule-Following). The results verify the presence of a circumplex structure in the value system.
@@
To evaluate the consistency and robustness of our multivariate system structure (i.e., the 5-dimensional relationship), we measure LLM values using two distinct prompt datasets: the psychometric prompts from GPV Ye et al. (2025b) and the red-teaming prompts from SALAD-Bench Li et al. (2024b), which feature distinct prompt distributions. Their measurements yield an average intra-LLM correlation of 0.87; here we use intra-LLM correlations because the relative value hierarchy within an LLM is more important than their absolute measurements. This result indicates a high level of consistency in the value structure across prompt distributions. We also find a high correlation (0.73) between intra-LLM value consistency and LLM safety scores Li et al. (2024b). It suggests that LLMs with higher value consistency tend to be safer. Complete results are available in §˜B.4.
@@ 5.3 Comparing Value Systems
We follow the standard validation procedures of CFA Schwartz and Boehnke (2004) to evaluate the structure validity of different value systems. For our value system, we construct it using half of the measurement data and bootstrap the data to ensure its sufficiency (#data points $\geq 5\times$ #variables). The other half of the data is held out for CFA. For Schwartz’s value systems, we map the observed value variables (the atomic values in our system) to its four high-level values or ten low-level values, according to the semantic relevance Schwartz and Boehnke (2004) measured by an embedding model OpenAI (2024); all data is used for CFA. Table˜2 displays the CFA results. Our value system demonstrates a better fit for the data, capturing the underlying values behind LLM generations.
@@
We follow the setup in (Ye et al., 2025b, Section 5.2) to predict LLM safety based on their value orientations. Table˜3 presents the prediction accuracy under different value systems. The higher accuracy of our value system indicates its superior predictive validity. In addition, according to the parameters of well-trained linear classifiers, we find that Social Responsibility, Rule-Following, and Rationality enhance safety, whereas Risk-taking and Self-Competence undermine it; see §˜B.6 for full results.
@@
We follow the experimental setup in (Yao et al., 2024a, Section 6.2) to perform LLM value alignment. Different value systems are respectively used to represent LLM outputs and desired human values. We employ GPV Ye et al. (2025b) as an open-vocabulary value evaluator for all value systems, but also include the original results of Yao et al. (2024a) using a Schwartz-specific evaluator for comparison. Table˜4 shows the alignment performance of different value systems. Alignment under our value system converges to the lowest harmlessness and the highest helpfulness, establishing its superior representation power. Full experimental details are available in §˜A.4.
@@ 6.2 Advantages of GPLA over Prior Psycho-Lexical Approaches
Here we discuss the main advantages of GPLA over the traditional psycho-lexical approach in constructing value systems. We defer the comparison with ValueLex Biedma et al. (2024), a recent study that constructs LLM value systems, to Appendix˜C.
@@
  • Comprehensive Coverage. LLMs can generate values in diverse linguistic forms, beyond words in dictionaries and thesauruses, and thus provide more comprehensive coverage of values. Please refer to §˜B.1 for the empirical evidence.
@@ Acknowledgements
This work is supported by the State Key Laboratory of General Artificial Intelligence; National Natural Science Foundation of China (Grant No. 62276006); Wuhan East Lake High-Tech Development Zone National Comprehensive Experimental Base for Governance of Intelligent Society; and the Fundamental Research Funds for the Central Universities.
@@ A.1 Data Statistics for Collecting Value Lexicons
We collect value-laden LLM generations from four data sources: ValueBench Ren et al. (2024), GPV Ye et al. (2025b), ValueLex Biedma et al. (2024), and BeaverTails Ji et al. (2024). They provide data of different forms: raw LLM responses, parsed perceptions (a sentence that is highly reflective of values Ye et al. (2025b)), and annotated values. The summary of the data statistics is shown in Table˜5.
@@ A.2 LLM Subjects
Our experiments involve 33 LLMs (Table˜6) and 21 profiling prompts, including the default prompt (’You are a helpful assistant’) Zhu et al. (2023) and 20 value-anchoring prompts Rozen et al. (2024).
@@ A.4 LLM Value Alignment
All experiments were conducted on two NVIDIA L20 GPUs, each with 48GB of memory. We generally follow the experimental setup in Yao et al. (2024a), with the exceptions noted below. As shown in Table˜4, our modifications improve the harmlessness of the aligned model with only marginal reduction in helpfulness.
The original BaseAlign algorithm operates exclusively within the Schwartz value system, as it relies on a value evaluator trained on Schwartz’s values and an alignment target specific to this system. We extend BaseAlign to align LLMs under any arbitrary value system. First, we employ GPV Ye et al. (2025b) as an open-vocabulary value evaluator. Second, we propose a method for distilling the alignment target from human preference data (§˜4.3). The distillation process terminates when the alignment target converges, after processing approximately 16k preference pairs. The results of this distillation are shown in Table˜7 for Schwartz’s values and Table˜8 for our system. The distilled alignment target, based on Schwartz’s values, closely matches the heuristically defined target in BaseAlign Yao et al. (2024a), demonstrating the effectiveness of our approach.
@@ B.2 Principal Component Analysis
We apply principal component analysis (PCA) to identify the underlying factors in our value system. The number of factors to retain is determined using both the scree plot and Cronbach’s alpha. The scree plot, shown in Fig.˜8, reveals an elbow at the fifth or sixth component. We retain five factors due to their high reliability.
@@ B.4 Consistency Across Datasets
In §˜5, we use the default profiling prompt for evaluating cross-dataset consistency. Table˜9 displays the intra-LLM correlation of the two value measurements for each LLM. In addition, we include the model safety scores from SALAD-Bench Li et al. (2024b) for reference. The results show that the intra-LLM correlation is generally high (0.87 on average), indicating the robustness of our value system. We also find that the intra-LLM value consistency is positively correlated with its safety (0.73).
@@ B.5 Correlation with Schwartz’s Values
We also measure the Schwartz’s values Schwartz (1992) of the LLMs subjects. Table˜10 presents the correlation between our factors and Schwartz’s values. The results can offer additional insights into our values.
@@ B.6 Contributions of Values to Safety
In the LLM safety prediction (§˜4.2) task, we train a linear classifier to predict the safety score of LLMs based on their values. The parameters of a well-trained classifier can be interpreted as the contributions of each value to the safety score. Such parameters, averaged over all trails, are gathered in Table˜12. The results show that Social Responsibility, Rule-Following, and Rationality positively contribute to safety, while Risk-taking and Self-Competence negatively contribute to safety.